# Practical techniques for high precision measurements on near-term quantum hardware: a Case Study in Molecular Energy Estimation

### v1

@@ Abstract
Achieving high-precision measurements on near-term quantum devices is critical for advancing quantum computing applications. In this paper, we explore several practical techniques to enhance measurement accuracy using randomized measurements, focusing on minimizing shot overhead, circuit overhead, measurement noise, and time-dependent measurement noise. Our approach leverages locally biased random measurements to reduce shot overhead, in addition to repeated settings and parallel quantum detector tomography to reduce circuit overhead and mitigate measurement noise. Additionally, we employ a blended scheduling technique to mitigate time-dependent measurement noise. We demonstrate the effectiveness of these techniques through a case study on the molecular energy estimation of the BODIPY molecule using the Hartree-Fock state on an IBM Eagle r3 computer, showcasing significant improvements in measurement precision. These strategies pave the way for more reliable and accurate quantum computations, particularly in applications requiring precise molecular energy calculations.
@@ I Introduction
In this paper we investigate how a combination of a set of strategies designed to address the primary issues associated with quantum measurements can impact the estimation of energy of a physical system. As a concrete example, we consider a molecular system, namely the Boron-dipyrromethene (BODIPY) molecule, which has been used in applications such as medical imaging, biolabelling, photoelectrochemistry, photocatalysis, artificial photosynthesis, optoelectronics, and photo-dynamic therapy. Simulations of such chemical compounds often aim to evaluate the molecular ground state energy, which can be done by preparing an ansatz state on the quantum computer using e.g. the Variational Quantum Eigensolver (VQE) algorithm [1, 2]. The estimation of certain molecular ground state energies on quantum computers has previously been demonstrated experimentally [1, 3, 4, 5, 6, 7, 8]. To obtain the energy of an ansatz state, one needs to evaluate the expectation value of the state and chemical Hamiltonian to a high precision. A commonly used upper bound for this precision, motivated by the sensitivity of reaction rates to changes in energy, is chemical precision [^1] For our measurement techniques, we use the term chemical precision instead of the more commonly used chemical accuracy, to distinguish between the precision in the estimation proceedure (e.g. the energy estimation of a state prepared by some algorithm, and the error in the estimation of a target energy of a molecule, e.g. the exact ground state energy of a molecule. at $1.6\times 10^{-3}$ Hartree [9]. The estimation of energies to chemical precision will be also a relevant challenge even in the age of fault-tolerant quantum computing due to the complexity in the observables that represent chemical molecules. In this work, we focus on decreasing the shot overhead (i.e. the number of times the quantum computer is measured), the circuit overhead (i.e. the number of times one needs to implement a different set of gates on the computer), as well as static and time-dependent measurement noise effects. Each of these issues poses significant obstacles to the accuracy and efficiency of quantum computations, and our strategies offer targeted solutions to mitigate their impact. We show that with the help of all of the aforementioned techniques, we are able to reach estimation errors close to chemical precision on current quantum hardware despite high readout errors on the order of $10^{-2}$ as well as the complexity of the observables.
@@
A key pillar of our measurement strategy is the use of informationally complete (IC) measurements. IC measurements offer several benefits such as allowing for the estimation of multiple observables from the same measurement data [10, 11, 12], a feature that has been proven useful for measurement-intensive algorithms such as ADAPT-VQE [13], qEOM [14] and SC-NEVPT2 [15], and providing a seamless interface between quantum and classical hardware, enabling the implementation of efficient error mitigation methods [16, 17]. But most importantly for the current study, IC measurements allow us to mitigate detector noise by performing quantum detector tomography [18, 19] and using the noisy measurement effects to build an unbiased estimator for the molecular energy. Furthermore, it also allows us to use efficient post-processing methods that enhance measurement accuracy and decrease shot overhead [20, 21, 22, 23]. Estimation of quantum chemical observables is also possible via non-IC measurements such as Pauli grouping [3, 24, 25, 26, 27, 28], which allows for the observable to be measured in groups of Pauli strings instead of individual Pauli strings on the quantum computer.
@@
Another technique we use to tackle shot overhead is the implementation of locally biased random measurements [29]. This technique allows us to choose measurement settings that have a bigger impact on the energy estimation, reducing the number of measurement shots required, while maintaining the informationally complete nature of the measurement strategy. Circuit overhead, another critical issue, is addressed through repeated settings and parallel Quantum Detector Tomography (QDT). These techniques optimize the use of quantum resources, allowing for more efficient circuit execution. Finally, we observe that the temporal variations of the detector might pose a barrier for achieving high precision measurements on current quantum hardware. To mitigate that, we introduce a blended scheduling technique to account for the dynamic noise.
@@
II Measurement strategies for high-precision measurements on near-term quantum computers
II.1 Unbiased estimators through informationally complete measurements
@@
where each projector corresponds to a projection on an eigenstate of the Pauli operators, and $p_{X}$, $p_{Y}$, and $p_{Z}$ correspond to the probabilities of picking each measurement direction. The measurement schemes we use in this work are particular instances of dilation-free IC-POVMs presented in Ref. [19]. In section II.5 we discuss how to improve the estimation precision by choosing the probability distribution of the local random measurements.
II.2 Repeated measurement settings for reducing circuit overhead
@@
In Appendix C we show that the estimator in Eq. (5) is unbiased and that the variance can be estimated as in Eq. (6):
@@
where $\langle\langle\omega\rangle_{i}^{2}\rangle$ denotes the conditional second moment, which is the second moment of the means for measurement setting $i$. In practice, this quantity can be estimated by keeping the results coming from different measurement settings separate, computing the mean squared for each and then taking its mean over all settings.
@@
A caveat with this method is that there is a fundamental limit in the precision that can be reached by only increasing the value of $T$, as the repeated settings scheme performs at best as well as the one-shot-per-setting scheme in terms of precision. We may investigate this limit using the inequality in Eq. (7). The derivation of the inequality is presented in Appendix C.
@@
II.3 Dealing with noisy detectors
@@
In our experiments we have used a technique called Quantum Detector Tomography (QDT) to obtain a accurate description of the noisy POVM effects. In particular, we have used the QDT method proposed in Ref. [35], which consists of measuring an informationally complete set of inputs (i.e. a set of states that span the Hilbert space such as the eigenstates of the Pauli matrices) and recovering the POVM effects through a likelihood method based on semi-definite programming. One of the main advantages of this method is that it can also incorporate errors in the set of input states.
@@
II.4 Dealing with temporal noise
@@
The fact that the detector is fluctuating in time is a serious threat to the measurement strategy based on IC-POVM and QDT we have introduced in Sec. II.1 and II.3. Indeed, if we run the circuits for QDT on real hardware at a different time with respect to the circuits for the estimation of the observable mean value as given by Eq. (1) and the detector on the quantum computer varies during this period of time, then the effects we are reconstructing through QDT do not correspond anymore to the ones we should use in Eq. (1), and the mean value of the observable may be significantly biased.
@@
We propose a new method to cope with temporal fluctuations that relies on a time average that gives rise to an unbiased estimator of the mean value of any observable. We call this procedure blending. With blending, one ensures that samples for all experiments are taken evenly with respect to chunks of time on the macroscopic level, i.e. throughout the entire experiment. This allows us to collect data for QDT on the same measurement process that is implemented for any other experiment run alongside it, even if the measurement process exhibits temporal fluctuations. A diagram depicting an abstract example of a fluctuating detector as well as a comparison between “regular” and blended scheduling is shown in Fig. 1. The method will produce an unbiased estimator even in the presence of temporal noise at the cost of a higher variance, which depends on the noise levels on the quantum computer. A mathematical demonstration of blending is shown in Appendix A and details about its implementation on IBM computers are provided in Appendix B.
@@
II.5 Decreasing shot overhead through biased randomised measurements
Specific points in the space of dilation-free POVMs can be identified as classical shadows [12] and locally biased classical shadows [29]. In the classical shadows measurement scheme, the measurement basis is chosen uniformly among the three non-trivial Pauli measurement operators $\{X,Y,Z\}$. The measurement in the basis of a Pauli operator has two possible outcomes, therefore a single-qubit effect of classical shadows on qubit $q$ is defined as
$\Pi_{k,i}^{(q)}=\frac{1}{3}E_{k,i}^{(q)},$ (8)
where $E_{k,i}^{(q)}$ is one of the eigenstates of the Pauli operator $P_{k}\in\{X,Y,Z\}$ that is measured on qubit $q$, determined by the measurement outcome $i$. An $N$-qubit classical shadows POVM is then defined as
$\Pi_{k,i}=\frac{1}{3^{N}}\bigotimes_{q=1}^{N}E_{k,i}^{(q)}.$ (9)
Locally biased classical shadows extend the classical shadows framework with non-uniform sampling probabilities [29]. The Pauli operators $\{X,Y,Z\}$ are locally sampled with a local probability $\tau_{k}^{(q)}$. All measurement outcomes $E_{k,i}^{(q)}$ arising from the Pauli operator $P_{k}^{(q)}$ have the same $\tau_{k}^{(q)}$. Thus, the measurement effects can be represented as the following
$\Pi_{k,i}=\bigotimes_{q=1}^{N}\tau_{k}^{(q)}E_{k,i}^{(q)}.$ (10)
For the purposes of our analysis, we implement locally biased classical shadows by determining the local probability distributions through the total weights of the qubit-marginalized Pauli operators in the input observable $O=\sum_{j}{c_{j}P_{j}}$. The sampling probability for the Pauli operator $P_{k}^{(q)}$ on qubit $q$ is determined by Eq. (11).
$\tau_{k}^{(q)}=\frac{\sum_{j}{|c_{j}|\mathbf{1}_{P_{k}^{(q)}=P_{j}^{(q)}}}}{\sum_{j}{|c_{j}|\mathbf{1}_{P_{j}^{(q)}\neq I}}},$ (11)
where $\mathbf{1}_{P_{k}^{(q)}=P_{i}^{(q)}}$ denotes the indicator function, giving $1$ if the Pauli operator $P_{k}^{(q)}$ equals the Pauli operator at qubit $q$ of the $j$th Pauli string in $O$ and 00 if not. The denominator is the absolute sum of all the coefficients corresponding to the Pauli strings in the observable where the identity operator is not found (or equivalently any one of $\{X,Y,Z\}$ are found) at qubit $q$. To guarantee informational completeness, we also use a heuristically chosen minimum threshold value $\tau_{\textrm{min}}=0.2$ for the values $\tau_{k}^{(q)}$, such that each Pauli operator is sampled locally with a non-zero probability. This procedure has the advantage that the sampling probabilities are inspired by the observable in question. Local Pauli operators with large contributions to the mean are measured more often than others, which reduces these terms’ contribution to the variance, while preserving informational completeness of the POVM.
The local Pauli operators can be measured by applying different single-qubit gates before the canonical measurement in the $Z$-basis, that is natively implemented on IBM hardware. To measure the $Z$ operator, one can directly perform the native measurement. To measure the $X$ operator, one applies a Hadamard gate before the canonical measurement. Lastly, to measure the $Y$ operator, one applies an $S^{\dagger}$ gate followed by a Hadamard gate and the canonical measurement.
III Case study: energy estimation of the BODIPY-4 molecule
For our analysis we consider the Boron-dipyrromethene (BODIPY) molecule. BODIPY and its derivatives are an important class of organic fluorescent dyes that possess attractive chemical and physical features such as high modularity, strong absorption of the visible light, significant fluorescence quantum yields, and exceptional thermal and chemical/photochemical stabilities. Due to these favorable characteristics, the BODIPY compounds has found widespread applications in medical imaging, biolabelling, photoelectrochemistry, photocatalysis, artificial photosynthesis, optoelectronics, and photo-dynamic therapy. This work considers only a segment of a project carried out in parallel regarding the simulation of the BODIPY molecule on quantum computer and we refer the reader to Ref. [39] for details regarding the complete experiment.
In this work we study the measurement of an initialization state of a $\Delta$ADAPT-VQE [40] procedure for an in-solvent BODIPY-4 molecule in various active spaces of 4e4o (8 qubits), 6e6o (12 qubits), 8e8o (16 qubits), 10e10o (20 qubits), 12e12o (24 qubits) and 14e14o (28 qubits). For each active space, we evaluate the energies of three Hamiltonians, each of which has a ground state being the actual ground state ($S_{0}$), first excited single state ($S_{1}$) and first excited triplet state ($T_{1}$) of the BODIPY-4 molecule in its active space.
Qubits 8 12 16 20 24 28
Num. Pauli strings 361 1819 5785 14243 29693 55323
The initialization state is represented by the Hartree-Fock state. Both the Hamiltonians and the initialization states are prepared in the Jordan-Wigner mapping. For the purposes of this experiment, we consider only the initialization state and not a state obtained through ground state preparation algorithms to avoid two-qubit gate errors, such that we can focus our analysis mostly on the measurement. As a result, the state in this experimental setup is very simple, however we note that the Hamiltonians contain a large number of Pauli strings and have a complex structure, which makes the measurement of such Hamiltonians to chemical precision (i.e. $1.6\times 10^{-3}$ Hartree) not trivial. We also note that the efficiency of our methods is not affected by the complexity of the state.
All three Hamiltonians are measured on the initialization state in a blended way, meaning that we have three sets of Hamiltonian-circuit pairs which are executed alongside circuits for QDT in a blended way such that we can assume that each experiment is performed with the same average measurement. In addition, the goal of the $\Delta$ADAPT-VQE is to estimate gaps between the different energies, which requires the estimations of the energy to be as homogeneous as possible. While we cannot ensure that the overall noise caused in the circuits is the same for all, through blending we can ensure that any temporal fluctuations in the noise is contained evenly in all circuits. The implementation of blending on IBM computers as well as details regarding executions is described in Appendix B. To assess the quality of the Hamiltonian-inspired locally biased classical shadows, in practice we perform different measurements for the three Hamiltonians, which means that even if the state that is prepared is the same for the three cases, the measurement is different and consequently the experiment performed on the quantum computer is as well.
Due to the informational completeness of the measurement schemes that we employ, it is possible to use the data from all three experiments to estimate the expectation values of all the Hamiltonians as the state, that is prepared for each, is the same. We note that even though we prepare the same initialization state for all three Hamiltonians, the $\Delta$ADAPT-VQE algorithm, for which the measurement schemes we present will be applied to, will ultimately produce different circuits for the different Hamiltonians. In such a case, it is generally impossible to reuse data from the experiment of one Hamiltonian-circuit pair to estimate the expectation value of an unrelated Hamiltonian-circuit pair as the circuits do not match. Even if the circuits do not match, such experiments can still be run in a blended way and the POVM effects obtained through QDT can be reused for all experiments if the measurement bases are the same.
IV Experimental Results
IV.1 Reduction of estimation bias using QDT
Figure 2: Errors obtained for the estimation of the 8-qubit $S_{0}$ Hamiltonian on the initialization state run on ibm_cleveland over the time period of 31.07.2024-03.08.2024. The curves indicate energies estimated using POVM effects with and without QDT as a function of the number of shots, averaged over 10 repetitions. The error bars represent the spread of the absolute errors over the 10 repeated experiments with a 95% confidence interval. The absolute error is computed as the absolute difference to the noiseless value of the initialization energy. The standard error is computed using the expression for the variance for repeated settings in Eq. (6).
IV.2 Measurement overhead of the measurement schemes
To demonstrate the viability of the measurement techniques presented in this work, we show results of experiments performed for the Hamiltonians in all the system sizes. We have in total run twelve separate experiments on ibm_cleveland during 11.07.2024-15.07.2024, where we measure each of the three Hamiltonians on the initialization state in parallel for all six active spaces that we have considered using both the Classical Shadows (CS) and Locally Biased Classical Shadows (LBCS) measurement strategies. Each Hamiltonian was measured using $S=3\times 10^{6}$ measurement settings with $T=100$ shots each. Alongside the measurement of the three Hamiltonians, we also performed QDT in a blended way, where each of the total 341 jobs contained 3 repetitions of each QDT circuit of the total twelve QDT circuits with each instance being measured for 100 shots, summing to a total of $1.02\times 10^{5}$ shots per QDT circuit and $1.23\times 10^{6}$ shots for the full QDT procedure. One run for one set of Hamiltonians for one active space size and measurement strategy took around 1 hour. A comparison between the LBCS and CS measurement schemes for the measurement of the $S_{0}$ Hamiltonian initialization energy is shown in Figs. 3 and 4. The results shown in the figure are obtained with effects characterized using QDT.
We show that QDT reduces errors in the estimation of the energies significantly also for these results in Sec. D. We can see that both measurement schemes contain the true energy within approximately $1\sigma$ of their error bars, however the LBCS scheme maintains a smaller standard error for all problems in the different system sizes compared to the CS scheme, meaning that on average LBCS will produce estimations with smaller variance. Despite that, we notice that for both measurement strategies the standard errors, highlighted by the shaded area, increase significantly with the number of qubits. The large variance presents a major problem for high precision experiments on quantum computers and is caused by both the measurement scheme and the measurement noise levels on the quantum computer. The effect of the increased variance due to QDT is discussed further in Appendix D.
IV.3 Characterization of temporal noise with blending
Lastly, we present an example of temporal noise that we have observed in one of the experiments shown in Sec. IV.1 in more detail. The experimental setup is identical to the one described in the aforementioned section with the exception that we have also run an additional job with only the twelve QDT circuits and $1\times 10^{5}$ shots each, a total of $1.2\times 10^{6}$ shots. This additional job represents a QDT experiment performed without blending, as it is fully performed before and thus independently from the experiment for the initialization state. In the presence of temporal noise, we expect the noise to be different between the two time periods in which the non-blended QDT experiment and the initialization state experiment are performed and therefore find a mismatch between the measurement that is implemented in both. In total, the experiment runtime was around 50 minutes and was performed on 03.08.2024 on ibm_cleveland.
V Conclusions
The errors reported for the Hamiltonians beyond 16 qubits are very large for applications in quantum chemistry. We note that the locally biased random measurements used for the measurement were only a simple heuristic implementation and can be significantly improved as in Refs. [31, 29]. In addition, a significant overhead is caused by the noise in the measurement process at such large qubit numbers, which must be mitigated by QDT. The reduction of variance at such large qubit numbers must be studied by applying different measurement schemes and measurement noise suppression strategies. The precision and accuracy of these results may also be improved by post-processing methods such as dual optimization [20] and TN-ICE [23]. We also note the existence of other measurement error mitigation techniques such as T-REX [34] and M3 [33]. How these methods could be applied to our IC-POVM formalism and the effect of temporal noise on these methods should be studied with care.

### v2

# Practical Techniques for High-Precision Measurements on Near-Term Quantum Hardware and Applications in Molecular Energy Estimation

@@ Abstract
Achieving high-precision measurements on near-term quantum devices is critical for advancing quantum computing applications. Quantum computers suffer from high readout errors, making quantum simulations with high accuracy requirements particularly challenging. This paper implements practical techniques to reach accuracies essential for quantum chemistry by addressing key overheads and noise sources. Specifically, we leverage locally biased random measurements for reducing shot overhead, repeated settings with parallel quantum detector tomography for reducing circuit overhead and mitigating readout errors, and blended scheduling for mitigating time-dependent noise. We demonstrate these techniques via molecular energy estimation of the BODIPY molecule on a Hartree-Fock state on an IBM Eagle r3, obtaining a reduction in measurement errors by an order of magnitude from 1-5% to 0.16%. These strategies pave the way for more reliable quantum computations, particularly for applications requiring precise molecular energy calculations.
@@ I Introduction
In this paper we investigate how a combination of a set of strategies designed to address the primary issues associated with quantum measurements can impact the estimation of energy of a physical system. As a concrete example, we consider a molecular system, namely the Boron-dipyrromethene (BODIPY) molecule, which has been used in applications such as medical imaging, biolabelling, photoelectrochemistry, photocatalysis, artificial photosynthesis, optoelectronics, and photo-dynamic therapy. Simulations of such chemical compounds often aim to evaluate the molecular ground state energy, which can be done by preparing an ansatz state on the quantum computer using e.g. the Variational Quantum Eigensolver (VQE) algorithm [1, 2]. The estimation of certain molecular ground state energies on quantum computers has previously been demonstrated experimentally [1, 3, 4, 5, 6, 7, 8]. To obtain the energy of an ansatz state, one needs to evaluate the expectation value of the state and chemical Hamiltonian to a high precision. A commonly used upper bound for this precision, motivated by the sensitivity of reaction rates to changes in energy, is chemical precision at $1.6\times 10^{-3}$ Hartree [9]. For our measurement techniques, we use the term chemical precision instead of the more commonly used chemical accuracy, to distinguish between the statistical precision in the estimation procedure (e.g. the estimation of an ansatz state energy) and the exact error of an ansatz state to a target energy of a molecule (e.g. the exact ground state energy of a molecule). The estimation of energies to chemical precision will be also a relevant challenge even in the age of fault-tolerant quantum computing due to the complexity in the observables that represent chemical molecules. In this work, we focus on decreasing the shot overhead (i.e. the number of times the quantum computer is measured), the circuit overhead (i.e. the number of times one needs to implement a different set of gates on the computer), as well as static and time-dependent measurement noise effects. Each of these issues poses significant obstacles to the accuracy and efficiency of quantum computations, and our strategies offer targeted solutions to mitigate their impact. We show that with the help of all of the aforementioned techniques, we are able to reach estimation errors close to chemical precision on current quantum hardware despite high readout errors on the order of $10^{-2}$ as well as the complexity of the observables.
@@
In this work we focus on the experimental implementation of informationally complete (IC) measurements. IC measurements offer several benefits such as allowing for the estimation of multiple observables from the same measurement data [10, 11, 12, 13, 14], a feature that has been proven useful for measurement-intensive algorithms such as ADAPT-VQE [15], qEOM [16] and SC-NEVPT2 [17], and providing a seamless interface between quantum and classical hardware, enabling the implementation of efficient error mitigation methods [18, 19]. But most importantly for the current study, IC measurements allow us to mitigate detector noise by performing quantum detector tomography [20, 21] and using the noisy measurement effects to build an unbiased estimator for the molecular energy. Furthermore, it also allows us to use efficient post-processing methods that enhance measurement accuracy and decrease shot overhead [22, 23, 24, 25]. Estimation of quantum chemical observables is also possible via non-IC measurements such as Pauli grouping [3, 26, 27, 28, 29, 30], which allows for the observable to be measured in groups of Pauli strings instead of individual Pauli strings on the quantum computer.
@@
Another technique we use to tackle shot overhead is the implementation of locally biased random measurements [31]. This technique allows us to choose measurement settings that have a bigger impact on the energy estimation, reducing the number of shots required, while maintaining the informationally complete nature of the measurement strategy. Circuit overhead, another critical issue, is addressed through repeated settings and parallel Quantum Detector Tomography (QDT). These techniques optimize the use of quantum resources, allowing for more efficient circuit execution. Finally, we observe that the temporal variations of the detector might pose a barrier for achieving high precision measurements on current quantum hardware. To mitigate that, we introduce a blended scheduling technique to account for the dynamic noise.
@@
II Results
II.1 Case study: Energy estimation of the BODIPY-4 molecule
For our analysis we consider the Boron-dipyrromethene (BODIPY) molecule. BODIPY and its derivatives are an important class of organic fluorescent dyes that possess attractive chemical and physical features such as high modularity, strong absorption of the visible light, significant fluorescence quantum yields, and exceptional thermal and chemical/photochemical stabilities. Due to these favorable characteristics, the BODIPY compounds has found widespread applications in medical imaging, biolabelling, photoelectrochemistry, photocatalysis, artificial photosynthesis, optoelectronics, and photo-dynamic therapy.
In this work we study the measurement of an initialization state of a $\Delta$ADAPT-VQE [32] procedure for an in-solvent BODIPY-4 molecule in various active spaces of 4e4o (8 qubits), 6e6o (12 qubits), 8e8o (16 qubits), 10e10o (20 qubits), 12e12o (24 qubits) and 14e14o (28 qubits). We provide an estimate of the energy of the ground state ($S_{0}$), first excited single state ($S_{1}$) and first excited triplet state ($T_{1}$) of the BODIPY-4 molecule in each of the selected active spaces. To estimate the excited state energies, we first use the techniques presented in Ref. [32] to generate Hamiltonians for which the original excited states become ground states, and then use the Hartree-Fock states of these generated Hamiltonians.
Qubits 8 12 16 20 24 28
Num. Pauli strings 361 1819 5785 14243 29693 55323
The initialization state is represented by the Hartree-Fock state, which is a separable state and does not require any two-qubit gates to be prepared. This choice comes from the fact that our focus is on the measurement errors alone, so we avoid two-qubit gate errors. Notice however, that the Hamiltonians contain a large number of Pauli strings and have a complex structure, which makes the measurement of such Hamiltonians to chemical precision (i.e. $1.6\times 10^{-3}$ Hartree) not trivial even for the Hartree Fock state.
All three Hamiltonians are measured on the initialization state in a blended way, meaning that we have three sets of Hamiltonian-circuit pairs which are executed alongside circuits for QDT in a blended way such that we can assume that each experiment is performed with the same average measurement. In addition, the goal of the $\Delta$ADAPT-VQE is to estimate gaps between the different energies, which requires the estimations of the energy to be as homogeneous as possible. While we cannot ensure that the overall noise caused in the circuits is the same for all, through blending we can ensure that any temporal fluctuations in the noise is contained evenly in all circuits. The implementation of blending on IBM computers as well as details regarding executions is described in Section III of the Supplementary Information [33]. To assess the quality of the Hamiltonian-inspired locally biased classical shadows, in practice we perform different measurements for the three Hamiltonians, which means that even if the state that is prepared is the same for the three cases, the measurement is different and consequently the experiment performed on the quantum computer is as well.
II.2 Reduction of estimation bias using QDT
Figure 1: Errors obtained for the estimation of the 8-qubit $S_{0}$ Hamiltonian on the initialization state run on ibm_cleveland. The curves indicate energies estimated using POVM effects with and without QDT as a function of the number of shots, averaged over 10 repetitions. The error bars represent the spread of the absolute errors over the 10 repeated experiments with a 95% confidence interval. The absolute error is computed as the absolute difference to the noiseless value of the initialization energy. The standard error is computed using the expression for the variance for repeated settings in Eq. (6).
II.3 Measurement overhead of the measurement schemes
To demonstrate the viability of the measurement techniques presented in this work, we show results of experiments performed for the Hamiltonians in all the system sizes. We have in total run twelve separate experiments on ibm_cleveland, where we measure each of the three Hamiltonians on the initialization state in parallel for all six active spaces that we have considered using both the Classical Shadows (CS) and Locally Biased Classical Shadows (LBCS) measurement strategies. Each Hamiltonian was measured using $S=3\times 10^{4}$ measurement settings with $T=100$ shots each. Alongside the measurement of the three Hamiltonians, we also performed QDT in a blended way, where each of the total 341 jobs contained 3 repetitions of each QDT circuit of the total twelve QDT circuits with each instance being measured for 100 shots, summing to a total of $1.02\times 10^{5}$ shots per QDT circuit and $1.23\times 10^{6}$ shots for the full QDT procedure. One run for one set of Hamiltonians for one active space size and measurement strategy took around 1 hour. A comparison between the LBCS and CS measurement schemes for the measurement of the $S_{0}$ Hamiltonian initialization energy is shown in Figs. 2 and 3. The results shown in the figure are obtained with effects characterized using QDT. Additional results for the $S_{1}$ and $T_{1}$ Hamiltonians are shown in Section VII of the Supplementary Information [33].
We show that QDT reduces errors in the estimation of the energies significantly also for these results in Section VI of the Supplementary Information [33]. We can see that both measurement schemes contain the true energy within approximately $1\sigma$ of their error bars, however the LBCS scheme maintains a smaller standard error for all problems in the different system sizes compared to the CS scheme, meaning that on average LBCS will produce estimations with smaller variance. Despite that, we notice that for both measurement strategies the standard errors, highlighted by the shaded area, increase significantly with the number of qubits. The large variance presents a major problem for high precision experiments on quantum computers and is caused by both the measurement scheme and the measurement noise levels on the quantum computer. The effect of the increased variance due to QDT is discussed further in Section VI of the Supplementary Information [33].
II.4 Characterization of temporal noise with blending
Lastly, we present an example of temporal noise that we have observed in one of the experiments shown in Sec. II.2 in more detail. The experimental setup is identical to the one described in the aforementioned section with the exception that we have also run an additional job with only the twelve QDT circuits and $1\times 10^{5}$ shots each, a total of $1.2\times 10^{6}$ shots. This additional job represents a QDT experiment performed without blending, as it is fully performed before and thus independently from the experiment for the initialization state. In the presence of temporal noise, we expect the noise to be different between the two time periods in which the non-blended QDT experiment and the initialization state experiment are performed and therefore find a mismatch between the measurement that is implemented in both. In total, the experiment runtime was around 50 minutes and was performed on ibm_cleveland.
III Discussion
In addition, we have applied our methods to system sizes of up to 28 qubits to study the scaling of the noisy measurement process on real quantum hardware. We note that the implementation of the techniques presented in this work will have little to no overhead with the number of qubits and are scalable in nature. However, the errors reported for the Hamiltonians beyond 16 qubits are very large for applications in quantum chemistry. We see two main causes for this increase in variance: The estimator and the measurement errors. To mitigate the effect of the variance in the estimator, one can explore the use of different measurement schemes, by choosing different POVMs as proposed in Refs. [35, 31, 36]. We note that the locally biased random measurements used for the measurement were only a simple heuristic implementation and can most definitely be improved upon. The precision and accuracy of these results may also be improved by post-processing methods such as dual optimization [22] and TN-ICE [25]. These post-processing techniques are fully compatible with the techniques we use in this paper, given that we use an accurate description of the measurement operators. The measurement errors that we have observed on hardware cause a significant measurement overhead at large qubit numbers, which can make high-precision measurements at a large scale challenging to implement. On one front, the gradual improvement in quantum hardware and their measurement fidelities will alleviate this overhead. On another front, the reduction of measurement errors at such large qubit numbers could be studied by exploring different measurement noise suppression strategies. For instance, other measurement error mitigation techniques such as T-REX [37], M3 [38] and ones proposed for sparse probability distributions in Ref. [39] may be applied to the IC-POVM formalism in conjunction with the techniques we have used in this paper. In addition, techniques based on quantum optimal control [40], which tailor control pulses based on noise, may improve the quality of the basis rotations performed before the computational basis measurement and Pauli twirling of the native measurement performed in T-REX may reduce the effect of measurement errors and make it easier for quantum detector tomography to obtain accurate and precise measurement operators.
Another natural next step for this research would be to study the performance and robustness of these techniques on more complex states and circuits e.g. ones produced by $\Delta$ADAPT-VQE. The study of such complex circuits on hardware will incur a significant amount of noise beyond the measurement process, which will need to be dealt with in some form using noise suppression [40], noise mitigation techniques [41, 19] or error correction [42] among others. The findings in this work will also benefit experimental implementations of algorithms that require informationally complete data such as the previously mentioned ADAPT-VQE [15], qEOM [16] and SC-NEVPT2 [17]. The techniques may also improve the experimental estimation of properties of quantum states as has been studied using simulations in Ref. [12]
IV Methods
IV.1 Unbiased estimators through informationally complete measurements
@@
where each projector corresponds to a projection on an eigenstate of the Pauli operators, and $p_{X}$, $p_{Y}$, and $p_{Z}$ correspond to the probabilities of picking each measurement direction. The measurement schemes we use in this work are particular instances of dilation-free IC-POVMs presented in Ref. [21]. In Section I of the Supplementary Information [33] we discuss how to improve the estimation precision by choosing the probability distribution of the local random measurements based on an approach introduced in Ref. [31].
IV.2 Repeated measurement settings for reducing circuit overhead
@@
In Section V of the Supplementary Information [33] we show that the estimator in Eq. (5) is unbiased and that the variance can be estimated as in Eq. (6):
@@
where $\langle\omega^{n}\rangle$ denotes the $n$-th moment of the random variable $\omega$ taken over all possible values of $\omega_{i}$ as defined in Eq. (1). In practice, the quantity can be estimated from all obtained samples $\omega_{m_{i,j}}$ as defined in Eq. (5) by averaging over them. The quantity $\langle\langle\omega\rangle_{i}^{2}\rangle$ denotes the conditional second moment, which is the second moment of the means obtained for each measurement setting $i$. In practice, this quantity can be estimated by keeping the results coming from different measurement settings separate, computing the mean squared for each and then taking its mean over all settings.
@@
A caveat with this method is that there is a fundamental limit in the precision that can be reached by only increasing the value of $T$, as the repeated settings scheme performs at best as well as the one-shot-per-setting scheme in terms of precision. We may investigate this limit using the inequality in Eq. (7). The derivation of the inequality is presented in Section V of the Supplementary Information [33].
@@
IV.3 Dealing with noisy detectors
@@
In our experiments we have used a technique called Quantum Detector Tomography (QDT) to obtain an accurate description of the noisy POVM effects. In particular, we have used the QDT method proposed in Ref. [44], which consists of measuring an informationally complete set of inputs (i.e. a set of states that span the Hilbert space such as the eigenstates of the Pauli matrices) and recovering the POVM effects through a likelihood method based on semi-definite programming. One of the main advantages of this method is that it can also incorporate errors in the set of input states.
@@
IV.4 Dealing with temporal noise
@@
The fact that the detector is fluctuating in time is a serious threat to the measurement strategy based on IC-POVM and QDT we have introduced in Sec. IV.1 and IV.3. Indeed, if we run the circuits for QDT on real hardware at a different time with respect to the circuits for the estimation of the observable mean value as given by Eq. (1) and the detector on the quantum computer varies during this period of time, then the effects we are reconstructing through QDT do not correspond anymore to the ones we should use in Eq. (1), and the mean value of the observable may be significantly biased.
@@
We propose a new method to cope with temporal fluctuations that relies on a time average that gives rise to an unbiased estimator of the mean value of any observable. We call this procedure blending. With blending, one ensures that samples for all experiments are taken evenly with respect to chunks of time on the macroscopic level, i.e. throughout the entire experiment. This allows us to collect data for QDT on the same measurement process that is implemented for any other experiment run alongside it, even if the measurement process exhibits temporal fluctuations. A diagram depicting an abstract example of a fluctuating detector as well as a comparison between “regular” and blended scheduling is shown in Fig. 5. The method will produce an unbiased estimator even in the presence of temporal noise at the cost of a higher variance, which depends on the noise levels on the quantum computer. A mathematical demonstration of blending is shown in Section II of the Supplementary Information [33], details about its implementation on IBM computers are provided in Section III of the Supplementary Information [33]. We also present a noise model for a fluctuating detector using unbalanced random telegraph noise in Section IV of the Supplementary Information [33]. Unbalanced random telegraph noise is a sign of non-Markovianity on superconducting qubit systems and has been a topic of research in the theoretical and experimental studies in Refs. [47, 48, 49, 50, 51].
@@
Data Availability
The data that support the findings of this article are available on Zenodo [52].
Code Availability
The code that was used to obtain the data that support the findings in this article cannot be made publicly available because it contains commercially sensitive information.